OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 3x - y = -3 and passes through the point (4, -2).

OpenStudy (mertsj):

If it is parallel to 3x-y=-3, it must have the same slope. The slope is -a/b of -3/-1 or 3. So write the equation of a line with slope 3 that contains (4,-2): y-(-2)=3(x-4) y+2 =3x-12 y=3x-14